I am having trouble setting up the system for the following problem:
The sum of Tom's age and Vanessa's age is 72. In 9 years the sum of their ages will be twice Tom's present age.
I know the first equation for the first sentence is: t + j = 72. What about the second sentence?
Thank you!
~Gina

Answer by stanbon(75887)

The sum of Tom's age and Vanessa's age is 72.
In 9 years the sum of their ages will be twice Tom's present age.
I know the first equation for the first sentence is: t + j = 72. What about the second sentence?
-----------------
t + j = 72
(t+9)+(j+9) = 2t
---
Substitute for "t" and solve for "j":
==================
Cheers,
Stan H.